# Shoot from the HIP: Hessian Interatomic Potentials Without Derivatives

### v1

@@ 1 Introduction
Over the past decades, molecular simulation has become a cornerstone for many tasks in material discovery and molecular design. Beyond energy and forces, chemists, chemical engineers, materials scientists, and physicists frequently rely on Hessians to enable many critical workflows (Qu & Bowman, 2016; Huang et al., 2018). For example, second-order geometry optimization accelerates the determination of reliable equilibrium structures. Transition-state searches are needed to find reaction pathways and estimate barrier heights and transition rates. This determines whether a reaction is viable and how to tune selectivity, with high industrial relevance for the chemical and pharmaceutical industry. Vibrational analyses connect theory to experiment through infrared and Raman spectra and provide zero-point energies, which are essential for ranking isomers and estimating reaction free energies.
Despite their broad utility, Hessian calculations remain a significant computational bottleneck. For a molecule with $N$ atoms, the Hessian is a $3N\times 3N$ matrix, where each entry requires a mixed second derivative of the electronic energy with respect to two nuclear coordinates
$\mathbf{H}_{IJ}^{\alpha,\beta}\;=\;\frac{\partial^{2}E}{\partial\mathbf{R}_{I}^{\alpha}\,\partial\mathbf{R}_{J}^{\beta}}$ (1)
@@
Traditional approaches rely either on analytic second derivatives or on numerical differentiation of gradients, both of which have severe limitations (Jørgensen & Simons, 1983; Handy & Schaefer, 1984; Komornicki & Fitzgerald, 1993). Analytic second derivatives are available only for a limited set of methods, such as Hartree-Fock, DFT, and MP2; whereas, for higher-level correlated wavefunction methods such as CCSD or CASPT2, they are either not yet implemented in most computational packages or computationally prohibitive. Numerical approaches sidestep analytic derivations but can become infeasible even for medium-sized molecules, requiring $O((3N)^{2})$ gradient evaluations (Helgaker et al., 2000).
Most computational chemistry calculations are carried out with Kohn-Sham density functional theory (Kohn, 1965) (KS-DFT), due to its balance of computational cost and accuracy (Bursch et al., 2022). However, the $O(N^{4})$ scaling of hybrid KS-DFT with system size remains expensive, which has spurred the development of machine-learning interatomic potentials (MLIPs). MLIPs promise to deliver near-DFT accuracy at a fraction of the cost by incorporating symmetries, thereby drastically reducing the need for extensive training data (Schütt et al., 2018; Thomas et al., 2018; Gasteiger et al., 2021; Schütt et al., 2021; Batatia et al., 2022; Batzner et al., 2022). Recently, some MLIP models have also been adapted to compute Hessians via automatic differentiation (AD). These AD Hessians incur significant computational costs of $O(N^{2})$, which presents practical challenges during training and inference (Yuan et al., 2024; Gönnheimer et al., 2025; Rodriguez et al., 2025; Williams et al., 2025; Cui et al., 2025).
@@
    We introduce a loss function that targets the part of the Hessian subspace particularly important for molecular optimization problems.
@@ 2 Background
2.1 MLIPs and equivariant neural networks
Early works on machine learning interatomic potentials used local atomic environment descriptors in combination with linear regression (Thompson et al., 2015; Shapeev, 2016), Gaussian processes (Bartók et al., 2010), and simple neural networks (Behler & Parrinello, 2007). SchNet (Schütt et al., 2018) was the first work to introduce rotation-invariant graph neural networks to predict molecular properties like energies and forces. Next-generation invariant architectures include ViSNet (Wang et al., 2022) and QuinNet (Wang et al., 2024). In parallel, rotation-equivariant architectures were developed, such as Cormorant (Anderson et al., 2019), DimeNet (Gasteiger et al., 2020), GemNet (Gasteiger et al., 2021), PaiNN (Schütt et al., 2021), NequIP (Batzner et al., 2022), SphereNet (Liu et al., 2022), MACE (Batatia et al., 2022) and the Equiformer family (Liao & Smidt, 2023; Liao et al., 2024). Graph neural networks represent a molecule as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in 3D Euclidean space $\mathbb{R}^{3}$, with nodes/atoms $I\in\mathcal{V}$ at positions $\mathbf{r}_{I}\in\mathbb{R}^{3}$ and edges $(I,J)\in\mathcal{E}$ defined by a cutoff radius. At layer $t$, each node carries a feature $\mathbf{h}_{I}^{(t)}$ that is a direct sum of irreducible $\mathrm{SO}(3)$ representations (irreps): $\mathbf{h}_{I}^{(t)}=\bigoplus_{l=0}^{l_{\max}}\mathbf{h}_{I}^{(t,l)}$, where $\mathbf{h}_{I}^{(t,l)}=\{\mathbf{h}^{(t)}_{I,l,m}\}_{m=-l}^{l}$ has $2l+1$ components. To update the node’s features, the model sends messages between each connected node, which are then aggregated in a permutation invariant fashion (usually the sum, or attention-weighted-sum operation) and then added to the receiving node. In $\mathrm{SO}(3)$-equivariant GNNs, node features transform under a global rotation $\mathbf{Q}$ via Wigner $\mathbf{D}$-matrices:
@@ 2.1 MLIPs and equivariant neural networks
Two irreps $\mathbf{p}_{l_{1}}$ and $\mathbf{g}_{l_{2}}$ interact using the Clebsch-Gordan tensor product (Thomas et al., 2018)
@@
where $\mathbf{Y}_{l_{2}}$ is the spherical harmonic of degree $l_{2}$, $w_{l_{1},l_{2},l_{3}}:\mathbb{R}\rightarrow\mathbb{R}$ is a learned weighting function, and $\mathbf{f}_{l_{1}}$ a simple function that takes both node features in, for example concatenation in the case of EquiformerV2. A more efficient variation of equation 4 relying on the reduction of the $SO(3)$ tensor product to the $SO(2)$ tensor products was proposed by eSCN (Passaro & Zitnick, 2023) and adopted by subsequent works, such as EquiformerV2. Finally, the messages are pooled to update the node features, which can be done, for example, via a sum or using graph attention (Liao et al., 2024).
@@ 2.2 Traditional methods for calculating Hessians
The gold standard for computing nuclear Hessians is to derive them analytically. In Kohn-Sham DFT, this is achieved using the Coupled Perturbed Kohn–Sham (CPKS) method (Deglmann et al., 2002), which is employed to calculate the ground truth in the dataset and in the workflows of our experiments. If available, analytical Hessians provide the most efficient ab-initio way to obtain Hessians, but they still scale as $O(N^{5})$. See Appendix A.1.1 for details. A major downside of the CPKS approach to Hessians is its highly complex implementation and need for exchange-correlation-functional-specific derivations, which makes it not always available. An alternative is to use a finite-difference scheme, requiring only energies or forces. Finite differences also scale $O(N^{5})$ for hybrid DFT, although with more numerical noise and a much larger computational prefactor. We provide more details in A.1.2.
@@
MLIPs have gained widespread popularity for their ability to accurately approximate molecular energies and forces. The forces can either be directly predicted (Gasteiger et al., 2021; Liao et al., 2024), or calculated as the derivative of the MLIP energy using automatic differentiation. Previous work has shown that the Hessian can be calculated using AD to some success (Yang et al., 2024; Yuan et al., 2024; Gönnheimer et al., 2025; Rodriguez et al., 2025; Williams et al., 2025; Cui et al., 2025). Although AD Hessians are much cheaper to calculate compared to DFT, they still have some downsides. First, accurately modeling the energy does not guarantee low errors on Hessians (Yuan et al., 2024), but requires dedicated training (Zhao et al., 2025; Cui et al., 2025). Second, compared to energies and forces, Hessians are expensive: Given a probing vector $\mathbf{v}$, AD only lets us calculate $\mathbf{H}\mathbf{v}$. To calculate the full $3N\times 3N$ Hessian, we need $3N$ Hessian vector products (HVPs) with the unit vectors $\mathbf{v}=\mathbf{e}_{i},i=0,...,3N-1$, each yielding one column of the Hessian. Since the energy calculation with an MLIP is usually $O(N)$, the total cost of the Hessian calculation is $O(N^{2})$. While the HVPs can theoretically be parallelized, the $O(N^{2})$ memory footprint requires even medium-sized systems to be computed sequentially.
@@ 2.4 Molecular optimization with Hessians
Geometry optimization locates minima $\operatorname*{arg\,min}_{\mathbf{R}}E(\mathbf{R})$ of the potential energy surface (PES) starting from non-equilibrium molecular geometries. It underpins most workflows from global searches to high-throughput screening. If Hessians are available, second-order methods are favored due to their fast convergence guarantees (Nocedal & Wright, 2006; Doikov et al., 2023). RFO (Simons et al., 1983; Banerjee et al., 1985) is a commonly used second-order optimizer (it makes use of Hessian information to accelerate convergence) for molecular geometries (Hermes et al., 2022). An attractive property of RFO is that it can be used both for minimizing $E(\mathbf{R})$ as well as for finding saddle points, and it is robust to indefinite Hessians (Banerjee et al., 1985). As computing the Hessian is usually too expensive, a common practice is to maintain an approximate Hessian using the BFGS quasi-Newton scheme. We provide more details in A.1.3.
@@
To carry out frequency analysis, one must first remove the five or six redundant degrees of freedom, corresponding to the invariance of the energy under rotation and translation. This is done by mass-weighing the Hessian and performing an Eckart projection (Louck & Galbraith, 1976), which we describe in A.6. Then, the projected matrix is decomposed into its eigenvalues, with all positive eigenvalues signaling a minimum, and exactly $N$ negative eigenvalue signaling the presence of a order-$N$ transition state.
@@ 3 HIP Hessian Prediction
Any equivariant neural network (Duval et al., 2024) with features of at least $l=2$ can be equipped with a Hessian readout head. In this work, we pick the EquiformerV2 architecture as the backbone with four transformer layers (Liao et al., 2024). Each layer contains a layer norm, message passing with graph attention, another layer norm, and a feed-forward layer.
@@ 3.2 Hessian Prediction Head
[…]
As we are reusing the message passing machinery with an interaction cutoff radius, predicting Hessians scales initially $O(N)^{2}$ in memory and compute within the cutoff, and then reduces to $O(N)$ scaling for larger systems. This sparsity with distance assumption aligns with the sparsity of Hessians due to locality of electron interactions that is common in quantum chemistry and can be mathematically rigorously justified (Kussmann et al., 2015). After message passing we have to transform the atom pair features $\mathbf{h}_{I,J,l,m,c}$ into the corresponding Hessian sub-blocks $\mathbf{H}_{I,J}$. First, we project the pair feature irreps down to a single $\{\mathbf{\tilde{h}}_{I,J,l,m}\}_{l\in\{0,1,2\}}$ irreps feature (1x0e + 1x1e + 1x2e in e3nn notation (Geiger & Smidt, 2022)) using a linear layer $\mathbf{W}_{l,c}$
$\mathbf{\tilde{h}}_{I,J,l,m}$ $=\mathbf{W}_{l,c}\mathbf{h}_{I,J,l,m,c}$ (7)
We then expand $\mathbf{\tilde{h}}_{I,J,l,m}$ to an intermediate Hessian sub-block $\mathbf{H}^{\prime}_{I,J}$ using the tensor product expansion (Unke et al., 2021):
@@
where $\mathbf{C}^{l,m}_{l_{1},m_{1},l_{2},m_{2}}$ are the Clebsch-Gordon coefficients ensuring that the relation in equation 5 is enforced (Unke et al., 2021). Physically, this is a simple change of basis from the coupled to the uncoupled angular momentum basis (Edmonds, 1996). $\mathbf{H}^{\prime}_{I,J,m_{1},m_{2}}$ is a $3\times 3$ block, since $m_{1}$ and $m_{2}$ run from $l_{1}=-l_{2}=-1$ to $l_{1}=l_{2}=1$. As $\mathbf{C}^{l,m}_{l_{1}=1,m_{1},l_{2}=1,m_{2}}=0$ for all $l>2$, $\mathbf{h}_{I,J,l,m,c}$ only needs to contain irreps up to $l=2$ to build the $3\times 3$ Hessian sub-blocks. Intuitively, equation 8 is the inverse of the irreducible tensor decomposition, so it assembles a spherical tensor back from its irreducible components. Conversely, this also means we need all irreps up to $l=2$ as only including irreps up to $l=1$ would not be sufficient to express the $3\times 3$ tensors that decompose into $l=2$ irreps. Finally, we symmetrize the immediate Hessian to the final prediction with $\mathbf{H}=\mathbf{H}^{\prime}+\mathbf{H}^{\prime\top}$.

### v2

@@ 1 Introduction
Over the past decades, molecular simulation has become a cornerstone for many tasks in material discovery and molecular design. Beyond energy and forces, chemists, chemical engineers, materials scientists, and physicists frequently rely on Hessians to enable many critical workflows (qu2016; huang2018). For example, second-order geometry optimization accelerates the determination of reliable equilibrium structures. Transition-state searches are needed to find reaction pathways and estimate barrier heights and transition rates. This determines whether a reaction is viable and how to tune selectivity, with high industrial relevance for the chemical and pharmaceutical industry. Vibrational analysis connects theory to experiment through infrared and Raman spectra and provide zero-point energies, which are essential for ranking isomers and estimating reaction free energies.
Despite their broad utility, Hessian calculations remain a significant computational bottleneck. For a molecule with $N$ atoms, the Hessian is a real $3N\times 3N$ matrix, where each entry requires a mixed second derivative of the electronic energy with respect to two nuclear coordinates
$\mathbf{H}=\nabla^{2}E,\ \ \ \ \mathbf{H}_{IJ}^{\alpha,\beta}\;=\;\frac{\partial^{2}E}{\partial\mathbf{R}_{I}^{\alpha}\,\partial\mathbf{R}_{J}^{\beta}}$ (1)
@@
Traditional approaches rely either on analytic second derivatives or on numerical differentiation of gradients, both of which have severe limitations (jorgensen1983; handy1984; komornicki1993). Analytic second derivatives are available only for a limited set of methods, such as Hartree-Fock, DFT, and MP2; whereas, for higher-level correlated wavefunction methods such as CCSD or CASPT2, they are either not yet implemented in most computational packages or computationally prohibitive. Numerical approaches sidestep analytic derivations but can become infeasible even for medium-sized molecules, requiring $O((3N)^{2})$ gradient evaluations (helgakerMulticonfigurationalSelfConsistentField2000).
Most computational chemistry calculations are carried out with Kohn-Sham density functional theory (kohn1965) (KS-DFT), due to its balance of computational cost and accuracy (bursch2022). However, the $O(N^{4})$ scaling of hybrid KS-DFT with system size remains expensive, which has spurred the development of machine-learning interatomic potentials (MLIPs). MLIPs promise to deliver near-DFT accuracy at a fraction of the cost by incorporating symmetries, thereby drastically reducing the need for extensive training data (schutt2018; thomas2018; gasteiger2021; schutt2021; batatia2022; batzner2022). Recently, some MLIP models have also been adapted to compute Hessians via automatic differentiation (AD). These AD Hessians incur significant computational costs of $O(N^{2})$, which presents practical challenges during training and inference (yuan2024; gonnheimer2025; rodriguez2025; williams2025; cui2025horm).
@@
    We show that downstream results can be improved further by using a loss function that targets the relevant Hessian subspace
@@ 2 Background
MLIPs and equivariant neural networks
Early works on machine learning interatomic potentials used local atomic environment descriptors in combination with linear regression (thompson2015; shapeev2016), Gaussian processes (bartok2010), and simple neural networks (behler2007). SchNet (schutt2018) was the first work to introduce rotation-invariant graph neural networks to predict molecular properties like energies and forces. Next-generation invariant architectures include ViSNet (wang2022) and QuinNet (wang2024). In parallel, rotation-equivariant architectures were developed, such as Cormorant (anderson2019), DimeNet (gasteiger2020), GemNet (gasteiger2021), PaiNN (schutt2021), NequIP (batzner2022), SphereNet (liu2022), MACE (batatia2022) and the Equiformer family (liao2023; liao2023equiformerv2). Graph neural networks represent a molecule as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ in 3D Euclidean space $\mathbb{R}^{3}$, with nodes/atoms $I\in\mathcal{V}$ at positions $\mathbf{r}_{I}\in\mathbb{R}^{3}$ and edges $(I,J)\in\mathcal{E}$ defined by a cutoff radius. At layer $t$, each node carries a feature $\mathbf{h}_{I}^{(t)}$ that is a direct sum of irreducible $\mathrm{SO}(3)$ representations (irreps): $\mathbf{h}_{I}^{(t)}=\bigoplus_{l=0}^{l_{\max}}\mathbf{h}_{I}^{(t,l)}$, where $\mathbf{h}_{I}^{(t,l)}=\{\mathbf{h}^{(t)}_{I,l,m}\}_{m=-l}^{l}$ has $2l+1$ components. To update the node’s features, the model sends messages between each connected node, which are then aggregated in a permutation invariant fashion (usually the sum, or attention-weighted-sum operation) and then added to the receiving node. In $\mathrm{SO}(3)$-equivariant GNNs, node features transform under a global rotation $\mathbf{Q}$ via Wigner $\mathbf{D}$-matrices:
@@
Two irreps $\mathbf{p}_{l_{1}}$ and $\mathbf{g}_{l_{2}}$ interact using the Clebsch-Gordan tensor product (thomas2018)
@@
where $\mathbf{Y}_{l_{2}}$ is the $2l_{2}+1$ dimensional vector of the spherical harmonics of degree $l_{2}$, $w_{l_{1},l_{2},l_{3}}:\mathbb{R}\rightarrow\mathbb{R}$ is a learned weighting function, and $\mathbf{f}_{l_{1}}$ a simple function that takes both node features in, for example concatenation in the case of EquiformerV2. A more efficient variation of equation 4 relying on the reduction of the $SO(3)$ tensor product to the $SO(2)$ tensor products was proposed by eSCN (eSCN2023) and adopted by subsequent works, such as EquiformerV2. Finally, the messages are pooled to update the node features, which can be done, for example, via a sum or using graph attention (liao2023equiformerv2).
@@
Equivariant matrix prediction
Similar to the Hessian, there are other matrices of interest that are rotation-equivariant and that previous work tried to directly regress. For example, PhisNet unke2021se, QHNet yu2023efficient, and WANet li2025enhancing try to predict the Fock matrix, and Graph2Mat febrer2025graph2mat predicts the density matrix.
@@ 2.1 Traditional methods for calculating Hessians
The gold standard for computing nuclear Hessians is to derive them analytically. In Kohn-Sham DFT, this is achieved using the Coupled Perturbed Kohn–Sham (CPKS) method (deglmannEfficientImplementationSecond2002), which is employed to calculate the ground truth in the dataset and in the workflows of our experiments. If available, analytical Hessians provide the most efficient ab-initio way to obtain Hessians, but they still scale as $O(N^{5})$. See Appendix A.1.1 for details. A major downside of the CPKS approach to Hessians is its highly complex implementation and need for exchange-correlation-functional-specific derivations, which makes it not always available. An alternative is to use a finite-difference scheme, requiring only energies or forces. Finite differences also scale $O(N^{5})$ for hybrid DFT, although with more numerical noise and a much larger computational prefactor. We provide more details in A.1.2.
@@
MLIPs have gained widespread popularity for their ability to accurately approximate molecular energies and forces. The forces can either be directly predicted (gasteiger2021; liao2023equiformerv2), or calculated as the derivative of the MLIP energy using automatic differentiation. Previous work has shown that the Hessian can be calculated using AD to some success (yang2024; yuan2024; gonnheimer2025; rodriguez2025; williams2025; cui2025horm). Although AD Hessians are much cheaper to calculate compared to DFT, they still have some downsides. First, accurately modeling the energy does not guarantee low errors on Hessians (yuan2024), but requires dedicated training (zhao2025; cui2025horm). Second, compared to energies and forces, Hessians are expensive: Given a probing vector $\mathbf{v}$, AD only lets us calculate $\mathbf{H}\mathbf{v}$. To calculate the full $3N\times 3N$ Hessian, we need $3N$ Hessian vector products (HVPs) with the unit vectors $\mathbf{v}=\mathbf{e}_{i},i=0,...,3N-1$, each yielding one column of the Hessian. Since the energy calculation with an MLIP is usually $O(N)$, the total cost of the Hessian calculation is $O(N^{2})$. While the HVPs can theoretically be parallelized, the $O(N^{2})$ memory footprint requires even medium-sized systems to be computed sequentially. rodriguez2025 in particular showed that including the Hessian in the loss leads to better extrapolation of energy and force predictions, less data for a given accuracy, and improved stability in MD. This comes at the cost of approximately 25 longer training times, which makes AD Hessians prohibitive to train.
@@ 2.3 Molecular optimization with Hessians
Geometry optimization locates minima $\operatorname*{arg\,min}_{\mathbf{R}}E(\mathbf{R})$ of the potential energy surface (PES) starting from non-equilibrium molecular geometries. It underpins most workflows from global searches to high-throughput screening. If Hessians are available, second-order methods are favored due to their fast convergence guarantees (nocedalNumericalOptimization2006; doikovSecondOrderOptimizationLazy2023). RFO (simons1983; banerjee1985) is a commonly used second-order optimizer (it makes use of Hessian information to accelerate convergence) for molecular geometries (sella2022). An attractive property of RFO is that it can be used both for minimizing $E(\mathbf{R})$ as well as for finding saddle points, and it is robust to indefinite Hessians (banerjee1985). As computing the Hessian is usually too expensive, a common practice is to maintain an approximate Hessian using the BFGS quasi-Newton scheme. We provide more details in A.1.3.
@@
To carry out frequency analysis, one must first remove the five or six redundant degrees of freedom, corresponding to the invariance of the energy under rotation and translation. This is done by mass-weighing the Hessian and performing an Eckart projection (louck1976), which we describe in A.6. Then, the projected matrix is decomposed into its eigenvalues, with all positive eigenvalues signaling a minimum, and exactly $N$ negative eigenvalue signaling the presence of a order-$N$ transition state.
@@ 3 HIP Hessian Prediction
Any equivariant neural network (duvalHitchhikerGuideGeometric2024a) with features of at least $l=2$ can be equipped with a Hessian readout head. In this work, we pick the EquiformerV2 architecture as the backbone with four transformer layers (liao2023equiformerv2). Each layer contains a layer norm, message passing with graph attention, another layer norm, and a feed-forward layer.
@@ 3.2 Hessian Prediction Head
$\mathbf{h}_{I,J,l,m,c}=\mathbf{v}_{I,J,l,m,c}=\mathbf{v}\big(\mathbf{h}^{(T)}_{I},\mathbf{h}^{(T)}_{J},\mathbf{r}_{I,J}\big)$ (6)
As we are reusing the message passing machinery with an interaction cutoff radius, predicting Hessians scales initially $O(N)^{2}$ in memory and compute within the cutoff, and then reduces to $O(N)$ scaling for larger systems. This sparsity with distance assumption aligns with the sparsity of Hessians due to locality of electron interactions that is common in quantum chemistry and can be mathematically rigorously justified (kussmann2015reduced). After message passing we have to transform the atom pair features $\mathbf{h}_{I,J,l,m,c}$ into the corresponding Hessian sub-blocks $\mathbf{H}_{I,J}$. First, we project the pair feature irreps down to a single $\{\mathbf{\tilde{h}}_{I,J,l,m}\}_{l\in\{0,1,2\}}$ irreps feature (1x0e + 1x1e + 1x2e in e3nn notation (geiger2022e3nn)) using a linear layer $\mathbf{W}_{l,c}$: $\mathbf{\tilde{h}}_{I,J,l,m}=\mathbf{W}_{l,c}\mathbf{h}_{I,J,l,m,c}$ We then expand $\mathbf{\tilde{h}}_{I,J,l,m}$ to an intermediate Hessian sub-block $\mathbf{H}^{\prime}_{I,J}$ using the tensor product expansion (unke2021se):
@@
where $\mathbf{C}^{l,m}_{l_{1},m_{1},l_{2},m_{2}}$ are the Clebsch-Gordon coefficients ensuring that the relation in equation 5 is enforced (unke2021se). Physically, this is a simple change of basis from the coupled to the uncoupled angular momentum basis (edmonds1996angular). $\mathbf{H}^{\prime}_{I,J,m_{1},m_{2}}$ is a $3\times 3$ block, since $m_{1}$ and $m_{2}$ run from $l_{1}=-l_{2}=-1$ to $l_{1}=l_{2}=1$. As $\mathbf{C}^{l,m}_{l_{1}=1,m_{1},l_{2}=1,m_{2}}=0$ for all $l>2$, $\mathbf{h}_{I,J,l,m,c}$ only needs to contain irreps up to $l=2$ to build the $3\times 3$ Hessian sub-blocks. Intuitively, equation 7 is the inverse of the irreducible tensor decomposition, so it assembles a spherical tensor back from its irreducible components. Conversely, this also means we need all irreps up to $l=2$ as only including irreps up to $l=1$ would not be sufficient to express the $3\times 3$ tensors that decompose into $l=2$ irreps. Finally, we symmetrize the immediate Hessian to the final prediction with $\mathbf{H}=\mathbf{H}^{\prime}+\mathbf{H}^{\prime\top}$.